OpenStudy (goformit100):

If ( x + 1, y – 2) = (3,1), find the values of x and y

OpenStudy (bahrom7893):

x=2, y = 3

OpenStudy (anonymous):

If (x+1,y-2) = (3,1), find the values of x and y So x + 1 = 3 and y - 2= 1 x = 2 and y = 3

OpenStudy (goformit100):

Thank you Madam.

OpenStudy (anonymous):

Yw